### v1

# Differentiable Voxelization and Mesh Morphing

Abstract.
In this paper, we propose the differentiable voxelization of 3D meshes via the winding number and solid angles. The proposed approach achieves fast, flexible, and accurate voxelization of 3D meshes, admitting the computation of gradients with respect to the input mesh and GPU acceleration. We further demonstrate the application of the proposed voxelization in mesh morphing, where the voxelized mesh is deformed by a neural network. The proposed method is evaluated on the ShapeNet dataset and achieves state-of-the-art performance in terms of both accuracy and efficiency. The code is available at https://github.com/Luo-Yihao/DOPH.
This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
1. Introduction
Figure 1. Differentiable voxelization algorithm converts a 3D mesh into a 3D grid of occupancy voxels by the winding number and solid angles on GPU. The first row shows hard cases of the voxelization of 3D meshes using traditional methods, where the mesh is not watertight, or containing non-regular triangles, or formed by combinations of shapes, or with complex geometry. The second row shows the voxelization results of the proposed differentiable voxelization algorithm followed by the re-cubifying.
Voxel is a fundamental representation of 3D objects in computer graphics and computer vision. It is widely used in 3D reconstruction (Brock et al., 2016; Wang et al., 2022), 3D rendering (Kato et al., 2020; Loper and Black, 2014; Liu et al., 2019), and 3D generative models (Mescheder et al., 2019; Guo et al., 2023; Ren et al., 2024) due to its uniform grid structure and simplicity. Therefore, voxelization, which converts a 3D mesh into a 3D grid of voxels, is an essential preprocessing step for many 3D deep learning tasks. However, the commonly used voxelization algorithms, such as the rasterization-based and the mesh-boolean-based methods (Wang, 2010; Zhou et al., 2016; Wang et al., 2022), implemented on the CPU, are not efficient and accurate, which comsumes a lot of time and memory and limits the development of high-resolution 3D deep learning works. To address this issue, we propose a novel differentiable voxelization algorithm for 3D meshes, which is fast by GPU acceleration, flexible in terms of abitrary smapling resolution on non-watertight meshes with abitrary topologies. More importantly, the proposed algorithm is differentiable, which enables the computation of gradients with respect to the input mesh, facilitating the optimization of mesh-level deep learning tasks. The proposed approach is based on the winding number and solid angles.
In recent work, Luo et al. (Luo et al., 2022) used the abstract inverse quadratic field to compute the occupancy (Mescheder et al., 2019) of voxels and slice a mesh differentiablely, which facilitates the slicing and voxelization of 3D meshes and achieves mesh reconstruction from 2D medical scan images. However, for general 3D meshes with non-regualr topologies and wiring structures, the abstract inverse quadratic field is not accurate enough due to the integration error. This issue is addressed by replacing field integration with the solid angles, which are more accurate and efficient in computing the occupancy of voxels without loss of differentiability. Jacobson et al. (Jacobson et al., 2013; Jacobson, 2016) achieved the voxelization and boolean operations of 3D meshes by the winding number and solid angles, but the naive computation of the solid angles leads a computational gap near surface boundaries, which requires further human interventions for discretization and smoothing. Further more, we discard the hierarchical winding number evaluation from (Jacobson et al., 2013) but implement the solid angles computation in a parallel manner on GPU, which is more efficient and accurate. Leveraging the proposed differentiable voxelization, we further demonstrate the application of the proposed voxelization in mesh morphing.
The main contributions of the differentiable voxelization algorithm proposed in this paper are as follows:
  • It is a fast and resolution-flexible algorithm for voxelization of 3D meshes, addmiting GPU-accelerated.
  • It is a differentiable algorithm for voxelization of 3D meshes, facilitating mesh morphing and other mesh-level optimization and deep learning tasks.
  • The proposed algorithm achieves state-of-the-art performance in terms of both accuracy and efficiency on the ShapeNet dataset.
2. Differentiable Voxelization
In this section, we introduce the technical details of the proposed differentiable voxelization algorithm for 3D meshes based on the winding number and solid angles.
2.1. Generalized Winding Number
The winding number is a fundamental concept in topology, which is used to determine the number of times a curve winds around a point. The winding number of a curve $\gamma$ around a point $p$ is defined as the number of times the curve winds around the point in the counterclockwise direction. The winding number is a topological invariant, which is invariant under continuous deformation of the curve. The winding number is defined as follows:
[…]
For 3D cases, the winding number of a surface $\Sigma$ around a point $p$ is defined as the number of times the surface winds around the point in the counterclockwise direction. The winding number is a topological invariant, which is invariant under continuous deformation of the surface. The winding number is defined as follows:
[…]
$\vec{E_{q}}$ $=\frac{\vec{r}-\vec{p}}{\left\|\vec{r}-\vec{p}\right\|^{3}}.$
For a closed boundary surface $\Sigma=\partial\Omega$, the winding number around the query point $p$ is equivalent the occupancy, which is a direct consequence of the Gaussian integral theorem (Maxwell, 1954), i.e.,
[…]
where $\delta(\vec{r}-\vec{p})$ is the Dirac delta function. The naive discretization of the winding number for a mesh $(V,F)$ can be formulated as follows: can be formulated as follows:
(4) $W(p,\Sigma)=$ $\frac{1}{4\pi}\sum_{i=1}^{F}\left\langle\vec{n}_{i},\vec{E_{q}}(c_{i})\right\rangle\Delta_{i},$
where $n_{i}$, $c_{i}$, and $\Delta_{i}$ are the normal vector, the center, and the area of the $i$-th face, respectively. When the mesh is well-regularized (e.g., triangular mesh constructed by the regular triangles) and the query point is far from the surface, the naive discretization of the winding number is accurate enough.
2.2. Solid Angles
However, general 3D meshes, especially the artifact meshes from manually designing and manufacturing, usually contain sharp edges, non-regular triangles, which leads to considerable errors in the naive discretization of the winding number. To address this issue, (Jacobson et al., 2013) proposed the hierarchical winding number evaluation by calulating solid angles, i.e.,
(5) $\text{Occp}(v,\Sigma)=W(p,\Sigma)=$ $\frac{1}{4\pi}\sum_{i=1}^{F}\Omega_{i}(q),$
where $\Omega_{i}(q)$ is the solid angle of the $i$-th face at the query point $q$. The solid angle (Van Oosterom and Strackee, 1983) is defined as the ratio of the area of the spherical cap to the square of the radius of the sphere, i.e.,
[…]
The mathematical equivalence between the solid angle and the winding number can be found in (Werner, 2017). For a triangle with vertices $v_{j}$, $j=0,1,2$, the solid angle can be computed as follows:
(7) $\Omega(q)$ $=\arctan\frac{\det(\vec{e}_{0},\vec{e}_{1},\vec{e}_{2})}{1+\left\langle\vec{e}_{0},\vec{e}_{1}\right\rangle+\left\langle\vec{e}_{1},\vec{e}_{2}\right\rangle+\left\langle\vec{e}_{2},\vec{e}_{0}\right\rangle},$
$\vec{e}_{j}$ $=\frac{\vec{v}_{j}-\vec{q}}{\left\|\vec{v}_{j}-\vec{q}\right\|}.$
Then the occupancy of a query point $q$ can be computed as the sum of the solid angles of all the faces from the mesh i.e., Eq. (5).
The naive computation of the solid angles like Eq. (5) leads a computational gap near surface boundaries due to the zero point of the denominator. (Luo et al., 2022) involves discretization and smoothing to achieve binary occupancy, which declines the accuracy. We replace $arctan$ by $\arctan 2$ to address this issue, where we prove that the $\arctan 2$ is continuous and smooth with zero gradients over query points coordinates, except the query points on the surface. Refer to the supplementary material for the proof. The proposed differentiable voxelization algorithm is presented by Algorithm 1.
Algorithm 1 Differentiable Voxelization
0:  Mesh $(V,F)$, query point $q$
0:  Occupancy ${\rm Occp}(q)$ Initialize ${\rm Occp}(q)=0$
  for each triangle $f_{i}$ in $F$ do
     $f_{i}=(v_{0},v_{1},v_{2})$, normalize $e_{i}=\frac{v_{i}-q}{\left\|v_{i}-q\right\|}$
     $\alpha=\det(e_{0},e_{1},e_{2}),\beta=1+\left\langle e_{0},e_{1}\right\rangle+\left\langle e_{1},e_{2}\right\rangle+\left\langle e_{2},e_{0}\right\rangle$
     ${\rm Occp}(q)={\rm Occp}(q)+\arctan 2\left(\alpha,\beta\right)$
  end for
  Return ${\rm Occp}(q)$
Fig. 1 shows an illustration of the differentiable voxelization of 3D meshes, including non-watertight meshes, meshes with non-regular triangles, and combinational and complex geometry, which causes the failure of traditional voxelization methods. The proposed differentiable voxelization algorithm achieves accurate and efficient voxelization of 3D meshes, which is essential for mesh morphing and other mesh-level optimization and deep learning tasks.
2.3. Mesh Morphing
Figure 2. Mesh morphing is a process of deforming a 3D mesh to another 3D mesh by a neural network. The proposed differentiable voxelization facilitates the optimization of mesh morphing, which admits the computation of gradients with respect to the input mesh under the supervision or conditions of voxelization ground truth.
Differentiable voxelization facilitates the optimization of mesh morphing, which admits the computation of gradients with respect to the input mesh under the supervision or conditions of voxelization ground truth. The voxelization ground truth can be usually obtained by CT scans (Buzug, 2011). Therefore the differentiable voxelization lays the foundation for 3D mesh reconstruction from 2D medical scan images (Litjens et al., 2017). It is worth noting that the differentiable voxelization is not limited to mesh morphing, but can be incorporated in other mesh-level optimization and deep learning tasks, such as mesh segmentation, mesh classification, and mesh generation.
Algorithm 1 provided an accurate differentiable voxelization algorithm for mesh morphing, which means the Eq. (5) converges to the ground truth of the voxelization. Even though Eq. (5) is composed by differentiable functions, the results approach the binary values of the voxelization ground truth. However, pros always come with cons. High accuracy of binary voxelization leads to gradients vanishing, despit the calculation is differentiable. In practice of mesh reconstruction of medical images, we use (4) to avoid the vanishing gradients, which is a trade-off between the accuracy and the differentiability. Fig 2 shows an illustration of the mesh morphing to fit the target segmentation of a human heart. Refer to (Luo et al., 2022) for more details of the cardiac mesh reconstruction via the differentiable voxelization.
3. Experiments
In the experiments, we evaluate our differentiable voxelization algorithm and other exiting methods on some of the meshes from (Krishnamurthy and Levoy, 1996) and (Jacobson, 2022). The methods evaluated includes the ”contains” method in Trimesh (Dawson-Haggerty et al., [n.d.]) library, the Mesh-to-SDF library (Kleineberg, 2024), and the SDF computing method used in (Wang et al., 2022), where we denoted as DOGN.
We evaluate the methods under different resolutions from 32 to 256, (32, 64, 128, 256). The Mesh-to-SDF method has two kinds of surface point methods namely ’scan’ and ’sample’. The ’Scan’ method, DOGN and our method are evaluated under all resolutions from 32 to 256. The ’Sample’ method and ’contains’ method are only evaluated under 32 and 64 which we think are sufficient to demonstrate their performance in terms of quality and time.
3.1. Settings and Details of the Methods
  • 1.
    Ours: since our method relies on GPU to process the data in parallel, we set the batch size of our method, i.e. the number of query points of each batch, to 2000, due to the memory limitation. Also, all the calculation is done under half-precision.
  • 2.
    Mesh-to-sdf (scan): For ’scan’ we used ’mesh-to-voxels’ API and set different resolutions directly.
  • 3.
    Mesh-to-sdf (sample): For the ’sample’ method, we set the batch size to 10000.
  • 4.
    Trimesh.contains: For trimesh’s contains method, we set the batch size to 10000.
  • 5.
    DOGN: No specific setting for DOGN, the level is set to 2/resolution as recommended.
All the meshes are normalized to [-1,1] before the computation of the ccupancy filed (SDF). All occupancy fields are firstly cubified and then smoothed with filter Laplacian method in Trimesh with 10 iterations and $\lambda=0.15$. Then, we compute the Chamfer distance and the Hausdorff distance with 20000 samples between the normalized original mesh and smoothed mesh 3 times, with the standard deviation reported. Results are computed under a single Nvidia RTX A6000 GPU and an AMD EPYC 7443P 24-core Processor.
3.2. Results
Name (Watertight) Num Faces Chamfer Distance Hausdorff Distance Time(s)
Diff ’Scan’ DOGN Diff ’Scan’ DOGN Diff ’Scan’ DOGN
Rocker-arm($\checkmark$) 20088 0.01702 0.05966 0.02648 0.02691 0.10362 0.03337 198.46 306.67 103.14
Beast 64618 0.01502 0.06272 0.02526 0.02309 0.10318 0.08098 744.10 368.34 107.00
Max Planck 99991 0.02033 0.37075 0.02555 0.18064 1.08849 0.03709 1158.30 484.46 135.59
Lucy($\checkmark$) 99970 0.01501 0.04233 0.02463 0.02503 0.11022 0.04462 1155.86 345.91 111.01
Cheburashka 13334 0.01806 0.06823 0.02769 0.02977 0.16839 0.08520 130.68 392.88 121.58
Happy 98601 0.01726 0.06299 0.02656 0.02687 0.15838 0.03588 1140.47 319.96 114.05
Teapot 6320 0.01757 0.25775 0.02353 0.10390 1.01302 0.03257 61.94 462.17 114.65
Bimba($\checkmark$) 224906 0.02056 0.09320 0.02989 0.03400 0.11556 0.03926 3050.98 455.75 135.22
Igea($\checkmark$) 268686 0.02164 0.13443 0.03212 0.04057 0.10788 0.03991 3649.88 534.83 135.75
Hourse($\checkmark$) 96966 0.01587 0.08751 0.02600 0.02425 0.60148 0.03771 1120.32 418.90 115.71
Homer($\checkmark$) 12000 0.01507 0.05253 0.02510 0.02412 0.09813 0.07579 117.49 341.66 107.74
Spot 5856 0.01996 0.08974 0.02957 0.03278 0.11833 0.03788 57.42 441.26 120.16
Armadillo($\checkmark$) 99976 0.01876 0.07367 0.02858 0.03146 0.11894 0.03571 1155.40 356.79 121.72
Suzanne 968 0.01947 0.09038 0.03012 0.08047 0.58337 0.09545 9.79 430.40 108.23
Stanford Bunny 69451 0.02215 0.20425 0.02847 0.10349 0.84211 0.04049 802.02 479.98 130.56
Nefertiti($\checkmark$) 99938 0.01779 0.08017 0.02747 0.03064 0.10384 0.03866 1156.84 560.41 118.79
Ogre 124008 0.01711 0.06440 0.02906 0.07934 0.13050 0.12526 1441.33 375.40 112.09
Cow($\checkmark$) 5804 0.01534 0.21486 0.02528 0.02932 0.95404 0.04360 57.02 354.76 105.84
Fandisk($\checkmark$) 12946 0.02054 0.08235 0.03002 0.03789 0.48765 0.04194 127.07 321.94 107.73
Car 157384 0.03552 0.07919 0.04584 0.16396 0.16026 0.18403 1849.66 285.01 129.18
Tableware 4788 0.01938 0.05631 0.20433 0.04357 0.17771 0.70841 47.18 247.94 102.09
Ladder 828 0.01530 0.06477 0.02357 0.02224 0.12011 0.03213 16.29 156.47 101.86
Bull 2418 0.01802 0.21368 0.02774 0.02930 0.81471 0.03632 24.04 373.57 112.22
Table 1. Results and Computation Times of Different Methods for resolution 256
It can be seen from table 1 that our method significantly outperforms ’Scan’ and DOGN in terms of quality (for results with standard deviation please refer to table 6 in Appendix A), under the resolution of 256. The cases in which the two distances have a large difference, the Max Planck for instance, is because the original mesh is not closed, and our method tends to fill those holes. The resulting mesh still captures the details of the original mesh and looks similar, as shown in figure give a car picture here ! . The time cost of our method is closely related to the number of faces of the mesh and the memory of the GPU, since our method relies on GPU memory, and computes the integral over all faces of a mesh. For example, our method processes mesh ’Ladder’ and ’Suzanne’ faster than the other two methods, but is significantly slower then process complex meshes like ’Igea’. However, with sufficient computing resources, our method will be significantly faster as shown in table 2. Otherwise, the processing time grows almost linearly to the number of faces and cubicly to the resolution when computing resources are limited.
For lower resolutions, we also include the ’contains’ and ’sample’ methods into consideration, our method produces comparable results to other methods as shown in table 2 and 3, 4.
Appendix A Appendix
Name (watertight) Num Faces Chamfer Distance(std) Time(s)
Diff ’Scan’ Trimesh ’Sample’ DOGN Diff ’Scan’ Trimesh ’Sample’ DOGN
Rocker-arm($\checkmark$) 20088 0.06632 (0.00014) 0.08379 (0.00013) 0.06656 (0.00013) 0.06641 (0.00008) 0.15448 (0.00035) 0.4251 12.3401 10.3270 72.9459 0.3631
Beast 64618 0.07167 (0.00011) 0.10516 (0.00032) 0.07189 (0.00019) 0.07183 (0.00034) 0.15194 (0.00068) 0.0075 14.3608 18.5567 72.6238 0.4087
Max Planck 99991 0.06057 (0.00013) 0.35081 (0.00080) 0.06569 (0.00022) 0.06266 (0.00009) 0.11371 (0.00066) 0.0065 16.9182 102.0209 76.7355 0.5564
Lucy($\checkmark$) 99970 0.05977 (0.00016) 0.07760 (0.00038) 0.05945 (0.00004) 0.05951 (0.00026) 0.15320 (0.00027) 0.0062 16.3613 66.0559 71.7929 0.4569
Cheburashka($\checkmark$) 13334 0.06254 (0.00012) 0.08655 (0.00028) 0.06275 (0.00019) 0.06270 (0.00024) 0.16430 (0.00055) 0.0045 12.6865 3.1635 72.2148 0.3245
Happy 98601 0.05548 (0.00013) 0.08143 (0.00032) 0.05548 (0.00021) 0.05537 (0.00007) 0.16028 (0.00066) 0.0073 17.0295 23.5146 71.7547 0.4803
Teapot 6320 0.05685 (0.00006) 0.12914 (0.00071) 0.05716 (0.00021) 0.06477 (0.00012) 0.16999 (0.00031) 0.0046 11.7145 4.4557 72.3538 0.3307
Bimba($\checkmark$) 224906 0.05728 (0.00014) 0.10217 (0.00025) 0.05718 (0.00021) 0.05718 (0.00013) 0.16595 (0.00026) 0.3337 20.9800 244.5197 75.5195 0.7512
Igea($\checkmark$) 268686 0.06211 (0.00006) 0.13745 (0.00031) 0.06203 (0.00016) 0.06215 (0.00003) 0.18204 (0.00030) 3.6248 24.0900 469.7210 78.1245 0.8369
Hourse($\checkmark$) 96966 0.06970 (0.00013) 0.12735 (0.00030) 0.07019 (0.00007) 0.10821 (0.00067) 0.15961 (0.00018) 0.0072 16.3146 87.9119 74.2368 0.4593
Homer($\checkmark$) 12000 0.05890 (0.00005) 0.07347 (0.00025) 0.05876 (0.00017) 0.05841 (0.00020) 0.15538 (0.00039) 0.0063 11.2691 2.3533 74.2110 0.3246
Spot 5856 0.06440 (0.00007) 0.10161 (0.00046) 0.06463 (0.00017) 0.06449 (0.00016) 0.16768 (0.00032) 0.0062 11.5103 9.6123 69.0767 0.3217
Armadillo($\checkmark$) 99976 0.07081 (0.00001) 0.11524 (0.00024) 0.07046 (0.00028) 0.07061 (0.00021) 0.16688 (0.00056) 0.0055 16.8934 121.1818 71.5409 0.5164
Suzanne 968 0.06440 (0.00024) 0.10288 (0.00031) 0.06312 (0.00003) 0.06173 (0.00017) 0.17534 (0.00040) 0.0064 12.1686 1.2807 69.1196 0.3132
Stanford Bunny 69451 0.06344 (0.00011) 0.15924 (0.00045) 0.06814 (0.00024) 0.06364 (0.00020) 0.12063 (0.00026) 0.0074 16.2307 81.7012 71.1485 0.5156
Nefertiti($\checkmark$) 99938 0.05991 (0.00011) 0.08721 (0.00039) 0.05972 (0.00022) 0.06000 (0.00023) 0.16238 (0.00036) 0.0055 17.6811 104.9907 75.9086 0.4892
Ogre 124008 0.06791 (0.00030) 0.09158 (0.00013) 0.06654 (0.00024) 0.06862 (0.00033) 0.15788 (0.00040) 0.0086 18.0386 12.8931 72.2672 0.5144
Cow($\checkmark$) 5804 0.06242 (0.00011) 0.21833 (0.00130) 0.06272 (0.00007) 0.17479 (0.00264) 0.15892 (0.00036) 0.0048 12.1376 1.2943 69.5902 0.2743
Fandisk($\checkmark$) 12946 0.05456 (0.00016) 0.09516 (0.00026) 0.05460 (0.00005) 0.05467 (0.00001) 0.16631 (0.00063) 0.0047 13.4979 6.9413 69.3316 0.3324
Car 157384 0.07015 (0.00014) 0.08850 (0.00011) 0.06765 (0.00013) 0.07037 (0.00034) 0.23775 (0.00040) 3.8466 20.1299 799.0143 76.3597 0.7136
Tableware 4788 0.63220 (0.00228) 0.21702 (0.00033) 0.62292 (0.00127) 0.62329 (0.00116) 0.17387 (0.00065) 0.0062 11.5668 3.6265 72.8735 0.3134
Ladder 828 0.28433 (0.00046) 0.29627 (0.00090) 0.28649 (0.00112) 0.28678 (0.00136) 0.13036 (0.00032) 0.1531 12.5374 0.08043 70.7119 0.2595
Bull 2418 0.06685 (0.00029) 0.21340 (0.00072) 0.06670 (0.00023) 0.13797 (0.00081) 0.16155 (0.00028) 0.0047 11.6457 2.0163 74.1523 0.3395
Table 2. Resulting Chamfer Distances and Computation Times of Different Methods for resolution 32
Name (watertight) Num Faces Chamfer Distance(std) Time(s)
Diff ’Scan’ Trimesh ’Sample’ DOGN Diff ’Scan’ Trimesh ’Sample’ DOGN
Rocker-arm($\checkmark$) 20088 0.03568 (0.00009) 0.06522 (0.00018) 0.03569 (0.00005) 0.03575 (0.00010) 0.08130 (0.00028) 3.10 15.80 97.47 470.83 1.90
Beast 64618 0.03357 (0.00002) 0.07294 (0.00014) 0.03351 (0.00006) 0.03761 (0.00013) 0.07956 (0.00020) 9.04 18.61 147.10 479.61 2.04
Max Planck 99991 0.03495 (0.00006) 0.35516 (0.00073) 0.03995 (0.00022) 0.03806 (0.00009) 0.05998 (0.00022) 14.30 22.87 806.17 500.46 2.56
Lucy($\checkmark$) 99970 0.03114 (0.00005) 0.05217 (0.00014) 0.03122 (0.00006) 0.03122 (0.00011) 0.07905 (0.00018) 14.07 19.51 574.55 479.88 1.99
Cheburashka($\checkmark$) 13334 0.03594 (0.00007) 0.07236 (0.00033) 0.03584 (0.00005) 0.03607 (0.00010) 0.08458 (0.00022) 1.46 16.67 26.07 474.34 2.16
Happy 98601 0.03270 (0.00014) 0.06755 (0.00029) 0.03266 (0.00004) 0.03296 (0.00007) 0.07941 (0.00042) 13.89 20.64 188.29 488.89 2.17
Teapot 6320 0.03291 (0.00014) 0.16490 (0.00093) 0.03879 (0.00005) 0.04875 (0.00032) 0.08601 (0.00013) 0.70 19.10 35.85 480.15 2.01
Bimba($\checkmark$) 224906 0.03427 (0.00007) 0.09563 (0.00037) 0.03432 (0.00007) 0.03445 (0.00012) 0.08473 (0.00016) 41.03 27.55 2199.34 497.14 2.74
Igea($\checkmark$) 268686 0.03663 (0.00006) 0.13448 (0.00028) * 0.03677 (0.00004) 0.09399 (0.00016) 52.59 31.82 $>$4000 507.74 2.86
Hourse($\checkmark$) 96966 0.03598 (0.00003) 0.09547 (0.00043) 0.03605 (0.00011) 0.06874 (0.00072) 0.08316 (0.00024) 15.53 22.91 791.05 484.71 2.14
Homer($\checkmark$) 12000 0.03298 (0.00012) 0.05691 (0.00013) 0.03292 (0.00010) 0.03292 (0.00007) 0.07907 (0.00023) 1.32 15.76 16.25 472.65 1.89
Spot 5856 0.03628 (0.00006) 0.09217 (0.00013) 0.03627 (0.00009) 0.03637 (0.00007) 0.08544 (0.00015) 0.65 19.74 81.19 478.09 2.05
Armadillo($\checkmark$) 99976 0.03737 (0.00005) 0.08675 (0.00014) 0.03718 (0.00003) 0.03747 (0.00005) 0.08546 (0.00012) 14.07 22.15 1001.82 478.16 2.32
Suzanne 968 0.03614 (0.00006) 0.09256 (0.00036) 0.03983 (0.00010) 0.03779 (0.00004) 0.08982 (0.00033) 0.11 17.65 11.48 474.98 1.87
Stanford Bunny 69451 0.03847 (0.00014) 0.16958 (0.00151) 0.04645 (0.00009) 0.03997 (0.00008) 0.06325 (0.00006) 9.76 22.77 669.95 475.62 2.50
Nefertiti($\checkmark$) 99938 0.03345 (0.00007) 0.07886 (0.00014) 0.03337 (0.00005) 0.03340 (0.00004) 0.08305 (0.00033) 14.07 24.16 908.54 503.51 2.30
Ogre 124008 0.03485 (0.00010) 0.07289 (0.00002) 0.03426 (0.00002) 0.03705 (0.00017) 0.08312 (0.00016) 19.13 21.47 126.22 481.03 2.20
Cow($\checkmark$) 5804 0.03376 (0.00008) 0.20966 (0.00108) 0.03378 (0.00014) 0.20101 (0.00081) 0.07951 (0.00016) 0.64 16.63 8.24 462.41 1.82
Fandisk($\checkmark$) 12946 0.04016 (0.00002) 0.08417 (0.00022) 0.04015 (0.00004) 0.04017 (0.00004) 0.07795 (0.00006) 1.42 16.30 61.16 462.34 1.86
Car 157384 0.04826 (0.00005) 0.08326 (0.00029) 0.04268 (0.00011) 0.05191 (0.00005) 0.14442 (0.00060) 28.38 23.19 2187.78 518.36 2.89
Tableware 4788 0.11128 (0.00041) 0.18683 (0.00079) 0.11143 (0.00022) 0.11268 (0.00042) 0.11983 (0.00033) 0.53 15.02 36.45 496.58 1.79
Ladder 828 0.13856 (0.00010) 0.23965 (0.00057) 0.13874 (0.00061) 0.13813 (0.00036) 0.07570 (0.00018) 0.39 14.08 0.25 482.66 1.60
Bull 2418 0.03542 (0.00003) 0.19739 (0.00125) 0.03541 (0.00003) 0.15641 (0.00064) 0.08279 (0.00019) 0.27 16.94 18.02 521.10 2.13
Table 3. Resulting Chamfer Distances and Computation Times of Different Methods for resolution 64
Name (watertight) Hausdorff Distance (resol 32) Hausdorff Distance (resol 64)
Diff ’Scan’ ’Sample’ ’Trimesh’ DOGN Diff ’Scan’ ’Sample’ ’Trimesh’ DOGN
Rocker-arm 0.09804 (0.00028) 0.13694 (0.00051) 0.09833 (0.00058) 0.09783 (0.00067) 0.21022 (0.00181) 0.04585 (0.00121) 0.11573 (0.00410) 0.04443 (0.00119) 0.04651 (0.00053) 0.11563 (0.00050)
Beast 0.19894 (0.00107) 0.26045 (0.00093) 0.19595 (0.00437) 0.20010 (0.00071) 0.23835 (0.00123) 0.08119 (0.00133) 0.13537 (0.00088) 0.07804 (0.00296) 0.45707 (0.00260) 0.11417 (0.00135)
Max Planck 0.17962 (0.00315) 1.09022 (0.00811) 0.27223 (0.00228) 0.22204 (0.00199) 0.21338 (0.00105) 0.17761 (0.00241) 1.07249 (0.00551) 0.26696 (0.00174) 0.63863 (0.00107) 0.10693 (0.00124)
Lucy 0.15998 (0.00132) 0.16109 (0.00193) 0.16288 (0.00137) 0.16278 (0.00154) 0.21806 (0.00112) 0.06911 (0.00104) 0.12530 (0.00233) 0.07138 (0.00029) 0.07018 (0.00013) 0.12193 (0.00040)
Cheburashka 0.11699 (0.00068) 0.15871 (0.00131) 0.11667 (0.00124) 0.11675 (0.00073) 0.22699 (0.00259) 0.08833 (0.00058) 0.13572 (0.00356) 0.08870 (0.00190) 0.17981 (0.00110) 0.14794 (0.00174)
Happy 0.12999 (0.00068) 0.14404 (0.00079) 0.12847 (0.00076) 0.12702 (0.00251) 0.23513 (0.00134) 0.06669 (0.00260) 0.12116 (0.00037) 0.07515 (0.00332) 0.13735 (0.00163) 0.11438 (0.00161)
Teapot 0.11404 (0.00119) 1.03774 (0.00159) 0.29660 (0.00184) 0.29425 (0.00041) 0.27024 (0.00118) 0.11140 (0.00194) 1.03101 (0.01343) 0.25777 (0.00379) 0.35074 (0.00100) 0.15234 (0.00113)
Bimba 0.10842 (0.00251) 0.17430 (0.00183) 0.10810 (0.00163) 0.13392 (0.00679) 0.27758 (0.00387) 0.05352 (0.00076) 0.14291 (0.00019) 0.05123 (0.00114) 0.12672 (0.05122) 0.13154 (0.00089)
Igea 0.08470 (0.00083) 0.16438 (0.00067) 0.08807 (0.00486) 0.08518 (0.00137) 0.22979 (0.00164) 0.04855 (0.00181) 0.12874 (0.00058) * 0.04498 (0.00081) 0.11116 (0.00069)
Horse 0.12941 (0.00116) 0.66342 (0.00346) 0.12695 (0.00019) 0.69284 (0.00014) 0.22085 (0.00111) 0.05391 (0.00248) 0.62432 (0.00143) 0.05171 (0.00058) 0.70203 (0.00186) 0.10906 (0.00110)
Homer 0.09541 (0.00109) 0.14960 (0.00024) 0.09477 (0.00110) 0.09332 (0.00086) 0.20811 (0.00181) 0.05552 (0.00177) 0.12214 (0.00030) 0.05250 (0.00038) 0.05271 (0.00154) 0.13304 (0.00102)
Spot 0.09370 (0.00125) 0.14076 (0.00066) 0.09496 (0.00147) 0.09580 (0.00093) 0.21575 (0.00071) 0.04495 (0.00060) 0.12255 (0.00149) 0.04177 (0.00042) 0.04207 (0.00109) 0.10402 (0.00070)
Armadillo 0.16884 (0.00294) 0.18805 (0.00160) 0.16510 (0.00644) 0.16510 (0.00147) 0.22296 (0.00237) 0.07307 (0.00485) 0.13891 (0.00158) 0.06400 (0.00390) 0.54865 (0.00217) 0.10959 (0.00144)
Suzanne 0.12759 (0.00192) 0.59997 (0.00491) 0.19467 (0.00321) 0.37024 (0.00129) 0.27440 (0.00019) 0.09585 (0.00327) 0.58307 (0.00278) 0.23128 (0.00116) 0.48175 (0.00127) 0.15865 (0.00179)
Stanford Bunny 0.10395 (0.00281) 0.70338 (0.00362) 0.21473 (0.00037) 0.23383 (0.00432) 0.20713 (0.00074) 0.09856 (0.00482) 0.84165 (0.01241) 0.23335 (0.00226) 0.36202 (0.00408) 0.10329 (0.00120)
Nefertiti 0.09277 (0.00073) 0.14624 (0.00248) 0.09496 (0.00052) 0.09448 (0.00147) 0.21426 (0.00053) 0.05853 (0.00330) 0.11916 (0.00113) 0.06024 (0.00078) 0.27661 (0.12798) 0.10759 (0.00192)
Ogre 0.15168 (0.00076) 0.16781 (0.00221) 0.15102 (0.00196) 0.15623 (0.00155) 0.21357 (0.00012) 0.09467 (0.00219) 0.12800 (0.00029) 0.06594 (0.00024) 0.16643 (0.00068) 0.15010 (0.00145)
Cow 0.17486 (0.00357) 1.02813 (0.00055) 0.17444 (0.00070) 1.05163 (0.00121) 0.21426 (0.00136) 0.08899 (0.00295) 0.96566 (0.00043) 0.09079 (0.00102) 0.99838 (0.00161) 0.10803 (0.00122)
Fandisk 0.10267 (0.00184) 0.12448 (0.00245) 0.09604 (0.00439) 0.10156 (0.00174) 0.20640 (0.00104) 0.06248 (0.00573) 0.22357 (0.00603) 0.05985 (0.00411) 0.06077 (0.00293) 0.11238 (0.00552)
Car 0.18345 (0.00453) 0.17601 (0.00163) 0.18431 (0.00303) 0.23542 (0.00072) 0.50652 (0.00330) 0.17009 (0.00174) 0.16606 (0.00402) 0.10175 (0.00323) 0.24027 (0.00020) 0.41532 (0.00032)
Tableware 0.89475 (0.00284) 0.44641 (0.00864) 0.86789 (0.00376) 0.86469 (0.00213) 0.64450 (0.00150) 0.21437 (0.00086) 0.38609 (0.00240) 0.22125 (0.00089) 0.41498 (0.00446) 0.67902 (0.00256)
Ladder 0.52241 (0.00009) 0.54227 (0.00024) 0.52273 (0.00013) 0.52240 (0.00060) 0.27069 (0.00046) 0.21872 (0.00060) 0.37963 (0.00067) 0.21957 (0.00044) 0.21828 (0.00047) 0.13241 (0.00110)
Bull 0.13289 (0.00373) 0.80594 (0.00121) 0.13217 (0.00162) 0.74641 (0.00350) 0.21890 (0.00068) 0.07061 (0.00363) 0.80143 (0.00205) 0.06779 (0.00362) 0.79966 (0.00378) 0.10765 (0.00041)
Table 4. Resulting Hausdorff Distances of Different Methods for resolutions 32 and 64
Name (watertight) Num Faces Chamfer Distance (std) Hausdorff Distance (std) Time(s)
Diff ’Scan’ DOGN Diff ’Scan’ DOGN Diff ’Scan’ DOGN
Rocker-arm($\checkmark$) 20088 0.02212 (0.00001) 0.06035 (0.00007) 0.04405 (0.00012) 0.02875 (0.00090) 0.10700 (0.00228) 0.05536 (0.00158) 24.91 48.49 14.67
Beast 64618 0.02020 (0.00003) 0.06433 (0.00011) 0.04267 (0.00005) 0.04024 (0.00282) 0.11064 (0.00161) 0.08628 (0.00031) 89.32 58.26 15.20
Max Planck 99991 0.02451 (0.00010) 0.36514 (0.00189) 0.03628 (0.00008) 0.17925 (0.00243) 1.07719 (0.00331) 0.05734 (0.00065) 139.89 75.70 18.59
Lucy($\checkmark$) 99970 0.01961 (0.00005) 0.04413 (0.00020) 0.04246 (0.00008) 0.03140 (0.00191) 0.11078 (0.00200) 0.07159 (0.00016) 139.63 52.06 13.55
Cheburashka($\checkmark$) 13334 0.02283 (0.00002) 0.06890 (0.00015) 0.04588 (0.00010) 0.03618 (0.00377) 0.12105 (0.00101) 0.10863 (0.00202) 15.71 60.38 16.87
Happy 98601 0.02165 (0.00007) 0.06306 (0.00022) 0.04383 (0.00008) 0.03353 (0.00252) 0.15536 (0.00016) 0.05918 (0.00079) 137.77 55.52 15.53
Teapot 6320 0.02224 (0.00002) 0.21243 (0.00083) 0.04506 (0.00018) 0.10466 (0.00050) 1.00711 (0.00192) 0.10269 (0.00028) 7.47 68.04 14.69
Bimba($\checkmark$) 224906 0.02426 (0.00006) 0.09357 (0.00016) 0.04720 (0.00009) 0.03871 (0.00186) 0.12011 (0.00015) 0.06437 (0.00091) 374.64 75.76 17.81
Igea($\checkmark$) 268686 0.02575 (0.00006) 0.13466 (0.00015) 0.05147 (0.00011) 0.03688 (0.00157) 0.11490 (0.00136) 0.06003 (0.00266) 451.48 87.85 18.27
Hourse($\checkmark$) 96966 0.02130 (0.00000) 0.08847 (0.00019) 0.04442 (0.00012) 0.03175 (0.00265) 0.60367 (0.00467) 0.05657 (0.00138) 135.42 66.81 14.92
Homer($\checkmark$) 12000 0.02014 (0.00001) 0.05297 (0.00021) 0.04231 (0.00009) 0.02928 (0.00239) 0.10570 (0.00048) 0.09825 (0.00138) 14.13 53.53 13.99
Spot 5856 0.02442 (0.00010) 0.09000 (0.00012) 0.04708 (0.00004) 0.03489 (0.00044) 0.11879 (0.00149) 0.05470 (0.00110) 6.93 66.23 14.88
Armadillo($\checkmark$) 99976 0.02365 (0.00009) 0.07687 (0.00006) 0.04697 (0.00010) 0.03633 (0.00140) 0.12065 (0.00033) 0.05840 (0.00132) 139.45 61.74 15.53
Suzanne 968 0.02423 (0.00001) 0.09001 (0.00010) 0.04917 (0.00018) 0.08368 (0.00175) 0.58039 (0.00234) 0.11331 (0.00174) 1.19 64.79 13.75
Stanford Bunny 69451 0.02668 (0.00004) 0.18646 (0.00057) 0.03927 (0.00010) 0.09665 (0.00154) 0.79551 (0.00374) 0.06049 (0.00240) 96.41 71.52 17.91
Nefertiti($\checkmark$) 99938 0.02221 (0.00003) 0.07914 (0.00014) 0.04485 (0.00004) 0.03048 (0.00077) 0.10623 (0.00101) 0.06128 (0.00267) 139.66 81.66 15.88
Ogre 124008 0.02241 (0.00004) 0.06616 (0.00023) 0.04644 (0.00016) 0.09947 (0.00153) 0.12944 (0.00218) 0.13369 (0.00087) 175.89 59.81 14.76
Cow($\checkmark$) 5804 0.02030 (0.00002) 0.21046 (0.00219) 0.04294 (0.00011) 0.03217 (0.00083) 0.95843 (0.00681) 0.05663 (0.00025) 6.88 52.27 13.42
Fandisk($\checkmark$) 12946 0.02753 (0.00005) 0.08065 (0.00017) 0.04490 (0.00013) 0.03944 (0.00156) 0.47152 (0.00420) 0.06074 (0.00087) 15.23 50.06 13.78
Car 157384 0.03895 (0.00022) 0.08014 (0.00016) 0.06685 (0.00001) 0.16614 (0.00562) 0.15815 (0.00457) 0.19055 (0.00186) 231.67 52.94 18.01
Tableware 4788 0.06436 (0.00009) 0.08177 (0.00029) 0.21082 (0.00053) 0.18588 (0.00271) 0.20589 (0.01163) 0.69894 (0.00030) 5.70 41.63 13.12
Ladder 828 0.02767 (0.00005) 0.08104 (0.00014) 0.04041 (0.00010) 0.04449 (0.00106) 0.14899 (0.00062) 0.06803 (0.00132) 2.40 30.80 13.08
Bull 2418 0.02309 (0.00007) 0.20242 (0.00165) 0.04534 (0.00008) 0.03539 (0.00165) 0.81664 (0.00358) 0.05535 (0.00029) 2.91 56.92 15.08
Table 5. Resulting Chamfer Distances, Hausdorff Distances, and Computation Times of Different Methods for resolution 128
Name (watertight) Num Faces Chamfer Distance (std) Hausdorff Distance (std) Time(s)
Diff ’Scan’ DOGN Diff ’Scan’ DOGN Diff ’Scan’ DOGN
Rocker-arm($\checkmark$) 20088 0.01702 (0.00008) 0.05966 (0.00015) 0.02648 (0.00006) 0.02691 (0.00079) 0.10362 (0.00084) 0.03337 (0.00053) 198.46 306.67 103.14
Beast 64618 0.01502 (0.00007) 0.06272 (0.00013) 0.02526 (0.00002) 0.02309 (0.00102) 0.10318 (0.00027) 0.08098 (0.00022) 744.10 368.34 107.00
Max Planck 99991 0.02033 (0.00002) 0.37075 (0.00201) 0.02555 (0.00002) 0.18064 (0.00128) 1.08849 (0.00818) 0.03709 (0.00180) 1158.30 484.46 135.59
Lucy($\checkmark$) 99970 0.01501 (0.00003) 0.04233 (0.00019) 0.02463 (0.00005) 0.02503 (0.00066) 0.11022 (0.00372) 0.04462 (0.00083) 1155.86 345.91 111.01
Cheburashka($\checkmark$) 13334 0.01806 (0.00003) 0.06823 (0.00002) 0.02769 (0.00012) 0.02977 (0.00083) 0.16839 (0.00756) 0.08520 (0.00111) 130.68 392.88 121.58
Happy 98601 0.01726 (0.00003) 0.06299 (0.00017) 0.02656 (0.00004) 0.02687 (0.00076) 0.15838 (0.00108) 0.03588 (0.00088) 1140.47 319.96 114.05
Teapot 6320 0.01757 (0.00007) 0.25775 (0.00125) 0.02353 (0.00005) 0.10390 (0.00108) 1.01302 (0.00292) 0.03257 (0.00040) 61.94 462.17 114.65
Bimba($\checkmark$) 224906 0.02056 (0.00006) 0.09320 (0.00020) 0.02989 (0.00010) 0.03400 (0.00125) 0.11556 (0.00130) 0.03926 (0.00236) 3050.98 455.75 135.22
Igea($\checkmark$) 268686 0.02164 (0.00006) 0.13443 (0.00021) 0.03212 (0.00002) 0.04057 (0.00378) 0.10788 (0.00041) 0.03991 (0.00023) 3649.88 534.83 135.75
Hourse($\checkmark$) 96966 0.01587 (0.00002) 0.08751 (0.00038) 0.02600 (0.00005) 0.02425 (0.00024) 0.60148 (0.00601) 0.03771 (0.00162) 1120.32 418.90 115.71
Homer($\checkmark$) 12000 0.01507 (0.00001) 0.05253 (0.00008) 0.02510 (0.00003) 0.02412 (0.00045) 0.09813 (0.00144) 0.07579 (0.00068) 117.49 341.66 107.74
Spot 5856 0.01996 (0.00005) 0.08974 (0.00039) 0.02957 (0.00008) 0.03278 (0.00201) 0.11833 (0.00163) 0.03788 (0.00242) 57.42 441.26 120.16
Armadillo($\checkmark$) 99976 0.01876 (0.00003) 0.07367 (0.00021) 0.02858 (0.00005) 0.03146 (0.00101) 0.11894 (0.00291) 0.03571 (0.00052) 1155.40 356.79 121.72
Suzanne 968 0.01947 (0.00003) 0.09038 (0.00011) 0.03012 (0.00013) 0.08047 (0.00191) 0.58337 (0.00194) 0.09545 (0.00056) 9.79 430.40 108.23
Stanford Bunny 69451 0.02215 (0.00004) 0.20425 (0.00038) 0.02847 (0.00014) 0.10349 (0.00072) 0.84211 (0.00584) 0.04049 (0.00102) 802.02 479.98 130.56
Nefertiti($\checkmark$) 99938 0.01779 (0.00001) 0.08017 (0.00040) 0.02747 (0.00006) 0.03064 (0.00078) 0.10384 (0.00123) 0.03866 (0.00489) 1156.84 560.41 118.79
Ogre 124008 0.01711 (0.00006) 0.06440 (0.00019) 0.02906 (0.00010) 0.07934 (0.01013) 0.13050 (0.00250) 0.12526 (0.00099) 1441.33 375.40 112.09
Cow($\checkmark$) 5804 0.01534 (0.00004) 0.21486 (0.00047) 0.02528 (0.00006) 0.02932 (0.00096) 0.95404 (0.00273) 0.04360 (0.00044) 57.02 354.76 105.84
Fandisk($\checkmark$) 12946 0.02054 (0.00007) 0.08235 (0.00028) 0.03002 (0.00006) 0.03789 (0.00251) 0.48765 (0.00281) 0.04194 (0.00159) 127.07 321.94 107.73
Car 157384 0.03552 (0.00008) 0.07919 (0.00037) 0.04584 (0.00010) 0.16396 (0.00319) 0.16026 (0.00826) 0.18403 (0.00240) 1849.66 285.01 129.18
Tableware 4788 0.01938 (0.00005) 0.05631 (0.00012) 0.20433 (0.00199) 0.04357 (0.00273) 0.17771 (0.00911) 0.70841 (0.00024) 47.18 247.94 102.09
Ladder 828 0.01530 (0.00004) 0.06477 (0.00031) 0.02357 (0.00009) 0.02224 (0.00056) 0.12011 (0.00055) 0.03213 (0.00047) 16.29 156.47 101.86
Bull 2418 0.01802 (0.00002) 0.21368 (0.00172) 0.02774 (0.00003) 0.02930 (0.00056) 0.81471 (0.00494) 0.03632 (0.00185) 24.04 373.57 112.22
Table 6. Resulting Chamfer Distances, Hausdorff Distances, and Computation Times of Different Methods for resolution 256